# Barrier-Riccati Synthesis for Nonlinear Safe Control with Expanded Region of Attraction

### v2

@@ III-C Local Safe Stability and ROA Expansion Condition
Nevertheless, local guarantees may not be sufficient and could be limited. For safety-critical systems, it is desirable to maximize the forward invariant region of the closed-loop system and similarly maximize the region of attraction (ROA) of the safe stabilization problem [Wang2018Permissive]. In what follows, using the regularity conditions needed for the SDRE approach, we derive a condition based on the problem matrices, namely the cost function design, when one can expand asymptotic stability to larger sets, potentially achieving semi-global asymptotic stability on an arbitrarily large set $\Omega_{c}$. Note that the first part of the proposition in the theorem only states the SDRE regularity assumptions.
@@
Consider the safety embedded optimal control problem (13) subject to the safety embedded system (11) with $\bar{A}(\bar{x}),\bar{g}(\bar{x}),Q(\bar{x})$, and $R(\bar{x})$ are at least $C^{1}$ and the pairs $\big(\bar{A}(\bar{x}),\bar{g}(\bar{x})\big)$ and $\big(\bar{A}(\bar{x}),Q^{\frac{1}{2}}(\bar{x})\big)$ are point-wise stabilizable and point-wise detectable respectively $\forall\bar{x}\in\bar{\Omega}$ such that the point-wise solution $P(\bar{x})$ to the BaS-SDRE (16) renders $\bar{A}_{c}(\bar{x})$ point-wise Hurwitz. The safety-critical system (2) under the safety embedded nonlinear feedback control law (14) is (semi-globally) asymptotically safely stable with ROA $\Omega_{c}=\{x\in\Omega,\ \bar{x}^{\top}P(\bar{x})\bar{x}\leq c\}$ for some $c>0$, if
@@
Then, by the hypothesis, $\dot{W}(\bar{x})\leq-\bar{x}^{\top}\hat{Q}(\bar{x})\bar{x}<0$ for all $\bar{x}\in\bar{\Omega}_{c}=\{\bar{x}\in\bar{\Omega},\ \bar{x}^{\top}P(\bar{x})\bar{x}\leq c\}$. Finally, by Lyapunov theory $\bar{\Omega}_{c}$ is positively invariant [khalil2002nonlinear], implying that $\bar{\Omega}_{c}$ is indeed an ROA for the closed-loop system. It remains to show how $\dot{P}$ can be explicitly computed.
Given that under the regularity assumptions the value function is smooth, $P(\bar{x})$ is differentiable. Hence, as in standard Lyapunov analyses, the time derivative $\dot{P}$ along the trajectory of the system of the solution $P$ of the SDRE can be determined by differentiating (16):
$\dot{P}(\bar{x})A_{c}(\bar{x})+A_{c}^{\top}(\bar{x})\dot{P}(\bar{x})+\tilde{Q}(\bar{x})=0$ (22)
@@
$\tilde{Q}(\bar{x})=P(\bar{x})\dot{A}(\bar{x})+\dot{A}^{\top}(\bar{x})P(\bar{x})-P(\bar{x})\dot{G}(\bar{x})P(\bar{x})+\dot{Q}(\bar{x})$
and the time derivatives of the elements of $A$, $G$ and $Q$ are computed as follows:
$\dot{A}_{ij}=\frac{\partial A_{ij}}{\partial\bar{x}}A_{c}(\bar{x})\bar{x},\ \dot{G}_{ij}=\frac{\partial G_{ij}}{\partial\bar{x}}A_{c}(\bar{x})\bar{x},\ \ \dot{Q}_{ij}=\frac{\partial Q_{ij}}{\partial\bar{x}}A_{c}(\bar{x})\bar{x}$
@@
$\text{vec}\big(\dot{P}(\bar{x})\big)=(A_{c}^{\top}(\bar{x})\otimes I+I\otimes A_{c}^{\top}(\bar{x}))^{-1}\text{vec}\big(\tilde{Q}(\bar{x})\big)$ (23)
@@
It is important to note that the state-feedback control law (14) derived via the SDRE is suboptimal in the sense that it enforces a quadratic structure on the value function and satisfies the HJB equation only point-wise. Consequently, while the controller guarantees asymptotic stability on $\Omega_{c}$, it does not, in general, solve the full nonlinear optimal control problem. Theorem 1, shows that $Q(\bar{x})$ is chosen to ensure that $Q(\bar{x})-\dot{P}\succ 0$ for all $\bar{x}\in\bar{\Omega}_{c}$ for the largest possible $c>0$, and consequently the largest possible ROA, $\bar{\Omega}_{c}$. It is worth noting that with an appropriate choice of $Q(\bar{x})$ (e.g., by fixing $R(\bar{x})$), $\bar{\Omega}_{c}$ can be made arbitrarily large, yielding a semi-global asymptotic stability of the origin of the safety embedded system.

### v3

@@ III-C Local Safe Stability and ROA Expansion Condition
Nevertheless, local guarantees may not be sufficient and could be limited. For safety-critical systems, it is desirable to maximize the forward invariant region of the closed-loop system and similarly maximize the region of attraction (ROA) of the safe stabilization problem [19]. In what follows, using the regularity conditions needed for the SDRE approach, we derive a condition based on the problem matrices, namely the cost function design, when one can expand asymptotic stability to larger sets, potentially achieving semi-global asymptotic stability on an arbitrarily large set $\Omega_{c}$. Note that the first part of the proposition in the theorem only states the SDRE regularity assumptions.
@@
Consider the safety embedded optimal control problem (13) subject to the safety embedded system (11) with $\bar{A}(\bar{x}),\bar{g}(\bar{x}),Q(\bar{x})$, and $R(\bar{x})$ are at least $C^{1}$, and the pairs $\big(\bar{A}(\bar{x}),\bar{g}(\bar{x})\big)$ and $\big(\bar{A}(\bar{x}),Q^{\frac{1}{2}}(\bar{x})\big)$ are point-wise stabilizable and point-wise detectable respectively $\forall\bar{x}\in\bar{\Omega}$ such that the point-wise solution $P(\bar{x})$ to the BaS-SDRE (16) renders $\bar{A}_{c}(\bar{x})$ point-wise Hurwitz. The safety-critical system (2) under the safety embedded nonlinear feedback control law (14) is (semi-globally) asymptotically safely stable with ROA $\Omega_{c}=\{x\in\Omega,\ \bar{x}^{\top}P(\bar{x})\bar{x}\leq c\}$ for some $c>0$, if
@@
Then, by the hypothesis, $\dot{W}(\bar{x})\leq-\bar{x}^{\top}\hat{Q}(\bar{x})\bar{x}<0$ for all $\bar{x}\in\bar{\Omega}_{c}=\{\bar{x}\in\bar{\Omega},\ \bar{x}^{\top}P(\bar{x})\bar{x}\leq c\}$. Finally, by Lyapunov theory $\bar{\Omega}_{c}$ is positively invariant [13], implying that $\bar{\Omega}_{c}$ is indeed an ROA for the closed-loop system. It remains to show how $\dot{P}$ can be explicitly computed.
Given that under the regularity assumptions, the value function is smooth, $P(\bar{x})$ is differentiable. Hence, as in standard Lyapunov analyses, the time derivative $\dot{P}$ along the trajectory of the system of the solution $P$ of the SDRE can be determined by differentiating (16):
$\dot{P}(\bar{x})\bar{A}_{c}(\bar{x})+\bar{A}_{c}^{\top}(\bar{x})\dot{P}(\bar{x})+\tilde{Q}(\bar{x})=0$ (21)
@@
$\tilde{Q}(\bar{x})=P(\bar{x})\dot{\bar{A}}(\bar{x})+\dot{\bar{A}}^{\top}(\bar{x})P(\bar{x})-P(\bar{x})\dot{\bar{G}}(\bar{x})P(\bar{x})+\dot{Q}(\bar{x})$
and the time derivatives of the elements of $\bar{A}$, $\bar{G}$ and $Q$ are computed as follows:
$\dot{\bar{A}}_{ij}=\frac{\partial\bar{A}_{ij}}{\partial\bar{x}}\bar{A}_{c}(\bar{x})\bar{x},\ \dot{\bar{G}}_{ij}=\frac{\partial\bar{G}_{ij}}{\partial\bar{x}}\bar{A}_{c}(\bar{x})\bar{x},\ \ \dot{Q}_{ij}=\frac{\partial Q_{ij}}{\partial\bar{x}}\bar{A}_{c}(\bar{x})\bar{x}$
@@
$\text{vec}\big(\dot{P}(\bar{x})\big)=-(\bar{A}_{c}^{\top}(\bar{x})\otimes I+I\otimes\bar{A}_{c}^{\top}(\bar{x}))^{-1}\text{vec}\big(\tilde{Q}(\bar{x})\big)$ (22)
@@
It is important to note that the state-feedback control law (14) derived via the SDRE is suboptimal in the sense that it enforces a quadratic structure on the value function and satisfies the HJB equation only point-wise. Consequently, while the controller guarantees asymptotic stability on $\Omega_{c}$, it does not, in general, solve the full nonlinear optimal control problem. Theorem 1 shows that $Q(\bar{x})$ is chosen to ensure that $Q(\bar{x})-\dot{P}\succ 0$ for all $\bar{x}\in\bar{\Omega}_{c}$ for the largest possible $c>0$, and consequently the largest possible ROA, $\bar{\Omega}_{c}$. It is worth noting that with an appropriate choice of $Q(\bar{x})$ (e.g., by fixing $R(\bar{x})$), $\bar{\Omega}_{c}$ can be made arbitrarily large, yielding a semi-global asymptotic stability of the origin of the safety embedded system.